Tucson Machinery, Inc., manufactures numerically controlled machines, which sell for an average price of $0.5 million each. Sales for these NCMs for the past two years were as follows: Use Exhibit 3.10.


LAST YEAR THIS YEAR
QUARTER QUANTITY (UNITS) QUARTER QUANTITY (UNITS)
I 9 I 12
II 15 II 21
III 23 III 25
IV 13 IV 17

a. Find the equation of a simple linear regression line using Excel. (Round your answers to 3 decimal places.)

Y =  +  t

b. Compute trend and seasonal factor from a linear regression line obtained with Excel. (Do not round intermediate calculations. Round your answers to 3 decimal places.)

Period Trend Forecast Seasonal Factors
Last Year I
II
III
IV
This Year I
II
III
IV

c. Forecast sales for next year using the above seasonal factors. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Period Forecast (Units)
Next Year I
II
III
IV
Answer #1

a)

Intercept of regression line = INTERCEPT(B2:B9,A2:A9) = 11.57143 = 11.571 (Rounding to 3 decimal place)

slope of regression line = SLOPE(B2:B9,A2:A9) = 1.178571 = 1.179(Rounding to 3 decimal place)

So, regression equation = 11.571 + 1.179x

b)

Period

Quantity(units)

Trend Forecast

Seasonal factors

1

9

12.75

0.706

2

15

13.929

1.077

3

23

15.108

1.522

4

13

16.287

0.798

5

12

17.466

0.687

6

21

18.645

1.126

7

25

19.824

1.261

8

17

21.003

0.809

Trend forecast is calculated by substituting x = 1,2,3,…8 into regression equation

Seasonal factors = Quantity/Trend forecast

c.

Next year Period

Cumulative period

Seasonal factors

Trend Forecast

Forecast(units)

1

9

0.696

22.182

15.45

2

10

1.102

23.361

25.73

3

11

1.392

24.54

34.15

4

12

0.804

25.719

20.67

We can see, Seasonal factor for period 1 next year = Seasonal factor for cumulative period 9 = average of seasonal factors of Period 1 of Last year and This year i.e average of cumulative period 1 and period 5 = average (0.706,0.687) = 0.696

Trend forecast for period 1 next year = Trend forecast for cumulative period 9 =11.571 + 1.179*9 = 22.182

Forecast for period 1 next year = forecast for cumulative period 9 = Trend forecast * seasonal factor for cumulative period 9 = 22.182*0.696 = 15.45

Likewise, forecast for other periods are calculated